40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1969     Entry time: Thu Nov 2 23:54:24 2017
Author: Craig 
Type: DailyProgress 
Category: TempCtrl 
Subject: Spooky Results from 2017 Oct 31 Beatnote Tracking 

Two days ago I set up the Agilent to keep taking spectra overnight while I tracked some relevant channels (elog 1966).  Today I plotted them the results.

Plot 1 shows the beatnote frequency [Hz] as a function of time [s].  Plot 2 shows some relevant channels over nearly the same timespan.  I have normalized the channels to the range [0,1] and put the min and max values of the channels in the legend.

There is some distinct correlation between the slow control voltages and the beatnote frequency.  The vaccan temperature stabilization seems to be working, even according to the out-of-loop sensor.

EDIT: Despite many attempts from several browsers and compressing the plot, I was unable to attach the second plot.  It's on the git repo here though if you care to look at it.

Attachment 1: 20171102_220215_TrackedBeatnote_Spectrum.pdf  156 kB  | Hide | Hide all
20171102_220215_TrackedBeatnote_Spectrum.pdf
Attachment 2: 20171102_220215_TrackedBeatnote.tgz  14 kB
ELOG V3.1.3-